# African Data Ethics: A Discursive Framework for Black Decolonial Data Science

### v1

@@ Appendix A PRISMA Diagram of Document Collection
A woman and a girl in white dresses sit in an open car.
@@ Appendix B African Data Ethics Framework Case Studies
The following section discusses five more case studies from data science work in the African context.
B.1. Decolonize & Challenge Internal Power Asymmetry Case Study: DRC Mineral Extraction
@@
The violent exploitation of miners in the Democratic Republic of Congo (DRC) is a harrowing example of how unchecked power from the West and within Africa corrupts the data science ecosystem. The DRC is home to an abundance of minerals necessary for data science. Specifically, cobalt and silicon are foundational components of all technology, especially the massive amount of computers that store and process data in global data centers (Ndjungu, 2020). To keep up with computing demand, multinational companies collect a copious amount of minerals from the DRC. Silicon and cobalt are often referred to as “blood minerals” because Western companies are able to make billions of dollars from the technology industry while the DRC continues to experience violent internal displacement, ineffective interventions, and a minuscule fraction of the value of their mining labor (Ndjungu, 2020). The extractive relationship between the Congolese and large Western companies has direct parallels to DRC’s colonial relationship with Belgium (Ndjungu, 2020). The monarch of Belgium, King Leopold, violently claimed the DRC to extract and sell raw materials so Belgium could be a major player in meeting the material demands of an industrializing Europe without any concern for the humanity of the Congolese people. This neocolonial relationship is furthered by African leaders who have assumed the role of middlemen in the mineral trade. The Dodd-Frank Section 1502 is a United States law passed to address the dehumanizing mining labor practices (Ndjungu, 2020). This law required companies to execute due diligence to ensure they were not selling DRC minerals mined from conflict. Rather than adhering to this law, multinational companies pulled out of direct agreements with the DRC and joined new partnerships with neighboring countries such as Rwanda. These other African countries serve as middlemen to buy blood minerals from the DRC and sell to the multinational companies so the companies could keep their hands clean in the eyes of the law (Ndjungu, 2020). Understanding the colonial and modern-day political background of DRC blood minerals is key to contextualizing calls for reducing the scale of datasets and demand for new technology 777https://newint.org/violence/2024/its-time-hold-big-tech-accountable-violence-drc. Every leader within the data science ecosystem has the responsibility to challenge and not perpetuate colonial and asymmetric power.
@@
B.2. Uphold Universal Good: International Partnership to Increase Access to COVID-19 Information
@@
B.3. Communalism in Practice Case Study: Challenging Utilitarian Data Ethics with a Communitarian Analysis
Another aspect of practicing communalism in data science is applying communitarian theories as a lens for evaluating data ethics. The Western concept of utilitarianism is a predominant paradigm in data ethics. Utilitarian data science aims to construct AI and other DDS that maximize the amount of social good and minimize the amount of social harm at scale (see effective altruists). African communitarian theories provide novel and strong critiques of utilitarian data ethics as well. One African data ethicist, in particular, applies African values to outline why utilitarianism: 1) trivializes human dignity through rationality, 2) justifies the suppression of non-dominant people and values, 3) ignores the role relationality plays in human-AI interaction, and 4) misinterprets the nature of self-sacrifice (Metz, 2021). Respectful debate with diverging perspectives is essential to the progress of RDS.
B.4. Invest in Data Institutions & Infrastructures Case Study: Building the Capacity of National Statistical Offices
B.5. Prioritize Education & Youth Case Study: African Data Science Research Mentorship
There is a growing community of African researchers, engineers, and technologists ready to utilize data science to make meaningful changes in their communities. However, the eagerness of young data science researchers must be met with mentorship and education on how to consider their positionality in their work. A set of fictional narratives inspired by real African data science projects illustrate the importance of education in the development of caring and successful data scientists (Abebe et al., 2021). Young researchers are interested in tackling large-scale issues to help their communities. However, this passion can lead to overstepping boundaries with data subjects, especially in cases of data refusal. Therefore, if a young data scientist is unsure, they should seek the advice of trusted advisors and inform themselves about the customs, languages, and history of the data communities they want to work with before collecting data. Even though within a data science team a young data scientist has lesser power due to seniority, they yield a lot of power over their data subjects. As such, it’s important to recognize the multiple positionalities a data scientist holds and adapt to feedback from all stakeholders.
@@ Appendix C Comparative Analysis Table
As a supplement to the Comparative Analysis with Particularist Data Ethics Frameworks, the following table shows the coding of minor African data ethics principles to each of the six particularist frameworks.
A table of minor African data ethics principles on the y-axis and six comparison papers on the x-axis ((Klein and D’Ignazio, 2024; Floridi et al., 2018; Carroll et al., 2021; Jobin et al., 2019; Younas and Zeng, 2024; Raquib et al., 2022). Between these axes white and green cell colors indicate if a principle is present or not. At the last row there are fractions to count each green cell (14/23, 14/23, 15/23, 13/23, 10/23, 6/23).

### v3

@@ Appendix A PRISMA Diagram of Document Collection
There is a note for the “identification of documents via databases and registers” that states: Databases included in search: Google Scholar, Web of Science, Scispace (LLM-driven literature search), ACM Digital Library and IEEExplore. The PRISMA flowchart starts with the identification of documents from databases and registers or other methods and documents of pruned due to exclusion criteria described in the methods data collection section until a final count of 47 documents are reached.
@@ Appendix B African Data Ethics Framework Case Studies
The following section presents six case studies from the African data ethics corpus to further contextualize our framework with real world examples from data science work in the African context.
B.1. Challenge Power Asymmetries Case Study: DRC Mineral Extraction
@@
The violent exploitation of miners in the Democratic Republic of Congo (DRC) is a harrowing example of how unchecked power from the West and within Africa corrupts the data science ecosystem. The DRC is home to an abundance of minerals necessary for data science. Specifically, cobalt and silicon are foundational components of all technology, especially the vast number of computers that store and process data in global data centers (Ndjungu, 2020). To keep up with computing demand, multinational companies collect a copious amount of minerals from the DRC with feigned ignorance of the unsafe working conditions, the prevalence of child labor, and the inhumane actions of the rebel groups who often control mining operations. Silicon and cobalt are often referred to as “blood minerals” because Western companies are able to make billions of dollars from the technology industry while the DRC continues to experience violent internal displacement, ineffective interventions, and a minuscule fraction of the value of their mining labor (Ndjungu, 2020). The extractive relationship between the Congolese and large Western companies has direct parallels to DRC’s colonial relationship with Belgium (Ndjungu, 2020). The monarch of Belgium, King Leopold, violently claimed DRC to extract and sell raw materials so Belgium could be a major player in meeting the material demands of an industrializing Europe without any concern for the humanity of the Congolese people. This colonial relationship is furthered by African leaders who have assumed the role of middlemen in the mineral trade. Dodd-Frank Section 1502 is a United States law passed to address dehumanizing mining labor practices (Ndjungu, 2020). This law required companies to execute due diligence measures to ensure they were not selling DRC minerals mined from conflict. Rather than adhering to this law, multinational companies pulled out of direct agreements with the DRC and joined new partnerships with neighboring countries such as Rwanda. These countries then acted as middlemen, buying conflict minerals from the DRC and selling them to multinational companies, allowing the companies to legally distance themselves from the exploitative sourcing(Ndjungu, 2020). Understanding the colonial and modern-day political background of DRC blood minerals is key to contextualizing calls for reducing the scale of datasets and demand for new technology 444https://newint.org/violence/2024/its-time-hold-big-tech-accountable-violence-drc. Every leader within the data science ecosystem has the responsibility to challenge and not perpetuate colonial and asymmetric power.
B.2. Assert Data Self-Determination Case Study: The Promise of African-led Data Science
Masakhane is an African natural language processing (NLP) collective that builds language datasets and models in Indigenous African languages. By all reports across documents, Masakhane practices all the principles proposed in this framework and especially upholds the major principle of Data Self-Determination (Eke et al., 2023c; Chan et al., 2021; Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023).They uphold these principles with a commitment to centering African values as their founding principles, working with existing public datasets to not infringe on data privacy, and building language datasets to preserve indigenous African languages for the future to come (Adelani et al., 2022). Masakhane also has a very welcoming and communal organizational structure that includes any interested party in weekly meetings and communications. Once a member wants to contribute to a Masakhane project, they must undergo in-house training to maintain quality of the their dataset and model. They also explicitly prohibit “parachute research from the Global North” to ensure the time and resources of their collective provide direct benefit to their communities.555https://www.masakhane.io/ Finally, every project plan includes a discussion of data privacy considerations to guide their work. While there are many other parts to Masakhane’s work, the practices described in their public documents demonstrate an African data science community that is closely aligned with the perspectives represented in our framework. Masakhane is part of the growing grassroots efforts to imagine and practice what African-led responsible data science can achieve.
B.3. Invest in Local Data Institutions & Infrastructures Case Study: Building the Capacity of National Statistical Offices
B.4. Utilize Communalist Practices Case Study: Challenging Utilitarian Data Ethics with a Communitarian Analysis
Another aspect of practicing communalism in data science is applying communitarian theories as a lens for evaluating data ethics. The Western concept of utilitarianism is a predominant paradigm in data ethics. Utilitarian data science aims to construct AI and other DDT that maximize the amount of social good and minimize the amount of social harm at scale (see effective altruists). African communitarian theories provide novel and strong critiques of utilitarian data ethics as well. One African data ethicist, in particular, applies African values to outline why utilitarianism: 1) trivializes human dignity through rationality, 2) justifies the suppression of non-dominant people and values, 3) ignores the role relationality plays in human-AI interaction, and 4) misinterprets the nature of self-sacrifice (Metz, 2021). Respectful debate with diverging perspectives is essential to the progress of responsible data science.
B.5. Center Communities on the Margins Case Study: Limits of Inclusive Representation in African Facial Recognition Technology
African people, and Black people in general, are severely underrepresented in facial recognition datasets, and this has led to a performance bias against Black users. African technology companies committed to addressing this bias so their primarily African users could rely on their products. For example, a women-led facial recognition start-up called BACE curated a diverse dataset from the local community so their system could better detect Black subjects (Eke et al., 2023c). Users upload photos of their IDs and short videos from their phones to confirm their identity 666https://www.bacegroup.com/. The technology was created to aid financial fraud investigation efforts in Ghana that are hampered by many citizens not having formal identification documents 777https://www.thehabarinetwork.com/meet-charlette-nguessan-she-and-her-team-have-innovated-facial-recognition-technology-designed-to-identify-black-africans. While at face value BACE is practicing community-engaged work, the principles described above call for more meaningful community involvement. Data representation is only one stage of the data science lifecycle that communities should be involved in. BACE was meeting the needs of financial institutions but did the team gauge if the local community was comfortable with this identification product (Birhane, 2020)? To center all communities in responsible data science, local communities should be involved in the ideation, creation, deployment, and maintenance of data-driven technology to build trust and buy-in for new technology (Okolo et al., 2023). The work of N’Guessan is very aligned with the principles of data self-determination & infrastructures because it was in indigenous effort to build a DDT for their local Black community. While this case study is seemingly doing everything “right” facial recognition is still used for surveillance and has performance issues for marginalized communities (Birhane, 2020).
@@ B.2. Assert Data Self-Determination Case Study: The Promise of African-led Data Science
B.6. Uphold Common Good: International Partnership to Increase Access to COVID-19 Information
@@ Appendix C Comparative Analysis Table
As a supplement to the discussion’s comparative analysis, the following table shows the coding of minor African data ethics principles to each of the seven particularist frameworks.
At the last row there are fractions to count each checked cell by data ethics framework (Data Feminism for AI: 10/19, AI4People: 11/19, CARE & FAIR Principles: 13/19, Global Landscape: 11/19, Central Asian AI ethics: 10/19, Islamic AI ethics: 6/19, Emancipatory Data Science: 10/19). In the last column, there are fractions to count each checked cell for each minor principle in an order consistent with the order of section 4 (3/7, 1/7, 3/7, 3/7, 4/7, 1/7, 5/7, 5/7, 6/7, 7/7, 5/7, 5/7, 0/7, 2/7, 1/7, 6/7, 3/7, 7/7, 4/7).